LHC 2010 - latest news

Week 14: Machine Coordinators:    Mike Lamont, Ralph Assmann

Latest plan: link (version of 6 April, 19:00)

List of End of Fill studies (in work)

Day Short summary/plan Link to reports from 8:30 and 17:00 meetings
Sunday 11th April 2010
  • Squeezed successfully in 1 and 5. Squeezed to 5.5 m in LHCb.
  • 01:53 dump after > 20 hours at flat top (SB & squeeze)
  • Pre-cycle Controls problems
  • 05:43 beam back
  • 10:00 – 16:00 Transverse feedback studies
  • 16:20 beam dump MKB vertical kicker (GEN D), PTU fault - access
  • Sunday evening: SPS RF problems postponed squeeze LHCb dipole switched off
  • Monday 01:25 Stable beams
  • 03:24 Lost cryo – S56 temperature sensor problem with ROD.A56B1, cryo-maintain back by 03:53
  • Ramp-down combo & precycle


Presentation at 08:30 meeting
Saturday 10th April 2010 Stable beams until 15:00, and then into adjust mode for squeeze commissioning. Presentation at 08:30 meeting
Friday 9th April 2010

Recovery from access during the morning, and start of a long session pushing up single bunch intensity at 450 GeV. Eventually succeeded in injecting 1.1 e11 (i.e. nominal bunch intensity), over-injecting onto a pilot.

A fill for physics was then established with stable beams being declared 06:15 Saturday morning.

Presentation at 08:30 meeting
Thursday 8th April 2010
  • Access
  • Evening: recovery/powering tests. Beam back overnight
Presentation at 08:30 meeting
Wednesday 7th April 2010
  • 00h00: Recovery. Access.
  • 07h00: Beam back.
  • 08h30: At 3.5 TeV.
  • 10:30: Stable beams
  • 12:53 Lost cryogenics maintain in MSL2 faulty valve on DFBM in 12 16:30 Beam back
  • 17:00 Ramp for squeeze QTF/QTD trips lost to RT trim following use of wrong tune references at start of ramp – will now systematically load references from sequencer
  • 20: 15 Ramp for squeeze
  • Overnight - successful squeeze to 2 m in Atlas and CMS
Presentation at the 08:30 meeting
Tuesday 6th April 2010
  • 01h00: Start of ramp.
  • 02h46: Stable beams.
  • 15h00: Beams lost due to undulator trip. Drift of the undulator RU.L4 resistive voltage signal during the last two days (no reset). During recovery RT trim checks.
  • 19h30: Beams at 450 GeV. b3 trims adjusted. Introduce b3 trim for change of chromaticity during snapback ("by hand").
  • 21:15: Start of ramp. Lost half of beam 1. No losses for beam 2. Chromaticity measured during the ramp for next iteration on b3.
  • 22:15: Beams lost at 3.5 TeV while preparing for the squeeze, due to RT trims. Reverted b3 trims.

Presentation at the 08:30 meeting

MP3 weekly summary

Monday 5th April 2010
  • Beams quietly colliding...
  • 13h30: Physics fill lost after 19 hours in physics mode! Lost cryo maintain in 78 for very short time.
  • 13h30-20h30: Recovery. Access required for RB.A87 switch problem. LHCb dipole polarity switched to positive.
  • 20h30-23h30: Pre-cycle and setting up for injection. A few MKI B2 issues.
  • Uptime: 17h. Unplanned maintenance: 7h. Technical stop: 0h. Physics: 13h30.
Presentation at the 09:00 meeting